40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 37 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  801   Thu Jan 26 12:40:21 2012 TaraNotesopticDual periscope base

 I fixed the drawing for periscope base. Will submit to the machine shop soon.

Attachment 1: dual_periscope_base.PDF
dual_periscope_base.PDF
  800   Thu Jan 26 00:34:23 2012 frank, taraDailyProgressBEATBeat improvement

 We investigated the beat noise, and found out a few issues we have to fix to improve the sensitivity.

  •  Periscopes

Resonant peaks from periscopes around 800 Hz are high. We use an aluminum bar to tighten the periscopes together. This bring down the peaks. However, the peaks are still high. We certainly have to work on the design for the periscope to get rid of the mechanical peaks.

2012_01_25_2.jpg

upper beam for dual periscope. This help reducing resonant peak around 800 Hz a lot.

 

  • Marconi's phase noise for ACAV loop.

The local oscillator we use for driving the AOM in ACAV loop is Marconi. We are sitting on its phase noise at 10kHz tuning range .When the tuning range is set to 1kHz, noise at frequency above 1kHz drops. However, with 1kHz tuning range setup, we don't have enough gain to suppress noise at lower frequency and noise around 100Hz goes up.

We need to characterize ACAV loop to project LO's phase noise on the noise budget (this has not been done yet). We might have to add another EOM for feedback on ACAV loop.

  •  Acoustic coupling

 We also noticed that talking noise can easily couple to the beat signal, but we are not certain where it happens. All mounts and posts will be checked if they are loosen or can be improved.

 

beat_20120125.png

  799   Wed Jan 25 23:41:38 2012 frank, taraDailyProgressopticnot much improvement on mode matching

We tried to redo the mode matching to RCAV by adjusting the lense position using translational stages. However the result does not improve that much. The visibility is still roughly the same at 96.5%. 

2012_01_25_3.jpg

 

We will do the mode match for ACAV. Right now the visibility for ACAV is ~90%.

We also monitored the beam reflected from ACAV. TEM02 shows up (see below figure), but we could not get rid of it by beam alignment. It is probably the distortion from the AOM.

2012_01_25_1.jpg

  798   Wed Jan 25 03:05:57 2012 frank, taraDailyProgressBEATBeat with 14.75 Mhz sideband

We added a ccd camera to monitor the reflected beam from the locked cavity. The shape of the beam(turns out to be LG10) tells us that we have to do a better mode matching.

==Motivation==

We found that bad alignment can cause extra noise in the beat noise, see  (psl:796).So we monitored the reflected beam from the locked cavity to check what kind of misalignment we have to fix. Currently, the shape of LG10 is the dominant one which tells us that we do not have a good spotsize/spot position match which can be fixed by moving the lenses in front of the cavity.

 2012_01_24_2.jpg

above, top right panel, the shape of the beam reflected from the locked RefCav.

2012_01_24.jpg

==To Do==

  So we will move the lenses in front of the cavity to optimize the mode matching.   The current value is ~ 96.5%, we want to improve it without spending too much time, not more than a day. We will mount the lenses on translational stages for better position adjustment.

  797   Wed Jan 25 02:51:44 2012 frank, taraDailyProgressopticFaraday Isolator is back

With the new EOM bases, we can place 2 EOMs and the Faraday Isolator back in to the setup.

The half wave plate(HWP) between the two EOMs is temporarily mounted with two posts mounted together on a cross holder, because there is not enough space. We will make a special post, so that it can be mounted between the EOMs.

 

2012_01_24_1.jpg

  796   Tue Jan 24 02:47:41 2012 frank, taraDailyProgressBEATBeat with 14.75 Mhz sideband

After switching to 14.75 MHz sideband setup, aside from the RFPDs, we did not change anything yet. We use TTFSS setup on the table. The seismic stack is still the same. The table is not floated. So we measure the beat  noise as a reference.

beat_2012_01_23.png

==Setup==

we use 14.75 resonant EOM for adding sideband, and the broadband EOM for feedback.  The resonant EOM is driven by an LO at 14.75 MHz @ 20dBm. The broadband EOM is supplied with 25V power supply, it is operated at low voltage because of the limited current source. 

RCAV is locked by TTFSS, ACAV is locked by UPDH, the phase shift for both loop is done by cable length adjustment.

 Note that we removed the Faraday isolator from the setup due to the limited space. It will be installed back later once we change the EOM base.

The power input is 0.2 mW on both cavities. The visibility is more than 80%.

==Problem and Plan==

We saw that beam alignment caused the bump around 1kHz to change significantly, so we will re-align everything to make sure that we have good mode matching.

The power input was chosen to be around 0.2 mW because at 1mW the error signal is irregular.  It might be the RFPD problem, we will look into it.

  795   Tue Jan 24 01:58:35 2012 FrankDailyProgressElectronics EquipmentRF-PDs

modified a second old PD for 14.75MHz. Had to replace many parts were not even close as indicated on the box with some labels nor the original values from the schematic. PD SN002 is slightly different to SN001 as i didn't have the parts to change it, e.g. the ac-gain is 11 instead of 10. Also had to replace the pd itself as it had a burn mark on it. PD is working fine, however not as good as the first one.

I also re-measured the noise vs photocurrent for the first one (SN001) with more averages but smaller span. Now they look really nice. RF-impedance is calculated using points with high photocurrent only.  Can fit to all data points later to get a slightly better number, but i think we really don't need to do as we don't care about a few Ohms.  Plots below:

SN_SN001.pngSN_SN002.png

 

  794   Mon Jan 23 11:36:18 2012 FrankDailyProgressRefCavrefcav work

Right, it's certainly good enough for now and the next months. However there are better opamps available which might give us a better SNR at lower power levels which i would like to try (and replacing opamps is only a few minutes). This would give us the chance to operate at lower power levels with better SNR, which might be useful later.

The error signal size mentioned has been measured with an un-characterized PD from the TNI, where i already mentioned that they are too small but i don't know what's inside the PD. We will check today what we get from the new PD and then either fix/modify the TNI PD or a second PD from LLO (where i have to replace the diode which has a big burn spot)

Quote:

 

 The PD seems certainly good enough - no need for new pre-amp. With such a high Q resonance, you might be limited by the 1 pA/rHz of the preamp, but if so that's OK.

Error signal size seems too small: for 1 mW and 6 kOhm from the RFPD...

 

  793   Mon Jan 23 09:16:43 2012 ranaDailyProgressRefCavrefcav work

 

 The PD seems certainly good enough - no need for new pre-amp. With such a high Q resonance, you might be limited by the 1 pA/rHz of the preamp, but if so that's OK.

Error signal size seems too small: for 1 mW and 6 kOhm from the RFPD...

  792   Mon Jan 23 01:49:30 2012 FrankSummaryElectronics EquipmentRFPD work

modified one of the photodiodes from LLO to be used as one of the RFPDs in our setup for locking the cavities. Changes are documented in the pdf (all files are on svn in data folder from 1/21/2012 and 1/22/2012).

  • Had to cut traces of the pcb to add a coil for the resonance circuit as there is no adjustable inductor with the right footprint with the large value we need for 14.75MHz (~3.5uH). Same for the notch which is not exactly at 29.5MHz, but adjusting that makes it impossible to adjust the main resonance exactly to 14.75MHz, limited by the range of the coil, as we don't have the right inductors. Coilcraft kits do not contain the values we need. Will order a couple of the right ones for future work.
  • Changed dc gain to 100, so DC output can be used up to 5mA but we are planning to use only ~1mW
  • RF gain is still 10, but might change that later after i checked the noise budget. Might replace the opamp and add a factor of 2 or 3 here if required.
  • PD dark noise is ~60nV/rtHz which is equivalent to 6.5e-12A/rtHz input noise, so shot noise limit equivalent of ~130uA of photocurrent. Confirmed with white light source.

SN_SN001.pngTF_SN001.png

RF-PD_SN001.pdf

Will continue modifying a second PD for the second cavity.

 

  791   Sat Jan 21 01:20:28 2012 FrankDailyProgressRefCavrefcav work

re-did the RF wiring for both cavities:

  • Power splitters are now located on the table below the TTFSS box. This keeps all cables from/to the PDs and EOMs very short
  • Power level at res. EOM is ~20dBm, LO is ~7dBm (TTFSS has an internal amplifier) - Master LO is still a SR345 with 23dBm.
  • Removed all old cables which we don't need anymore and moved the ones used from the table to the framing structure to get less seismic coupling to the floating table from them.
  • Finished phase adjustment for the TTFSS and locked the cavity using all three actuators.
  • Temp feedback to the laser is now done by the TTFSS unit. There is NO computer control for any of the FSS stuff anymore - only for the PMC. cable to laser unit is temporary as we don't have the right cable length available.
  • Power is adjusted to 1mW (total)
  • We still don't have enough gain. Common gain is at max, fast at 782, (offset 500).
  • Measured in-loop ep noise floor - will add plot later. Noise floor measured at OUT1 of TTFSS box (common path, mixer output) is ~10nV/rtHz
  • Error signals are still pretty tiny (230 mVpp)
  • re-wired beat setup. Need to add a coupler to hook up the frequency counter in parallel. Will use that for characterization of radiative heater and temp feedback to it for beat note stabilization.

Next steps:

  • Full characterization of 14.75MHz res. PD from TNI
  • build a second PD for 14.75MHz.
  • add some channels to DAQ - there are currently no FSS channels at all
  • measure beat note by Tue
  • finish in-vac wiring
  • open vacuum can and install rad shields+heater
  • while in air measure TF from table to top stack with small accelerometer and compare with model
  790   Fri Jan 20 15:31:08 2012 FrankDailyProgressopticRCAV is locked

Don't change the way the EOM is wired !! If you do so the case is not connected to GND/protective earth anymore and your high voltage is on the metal case!

Quote:

RCAV is locked using TTFSS.

I have not tried to measure noise at the error point yet, since I have to flip the phase by 180 for feedback to EOM (TTFSS has a phase flip switch for FAST feedback only). I used a long BNC cable to change the phase by 180 degree, I think making an adaptor for EOM connector to flip the signal might be a better idea to try.

 

  789   Thu Jan 19 23:43:43 2012 taraDailyProgressopticRCAV is locked

RCAV is locked using TTFSS.

It took awhile before I could lock the cavity because the 14.75 MHz EOM tilts the beam path, and I had to realign the beam. We don't know why the beam path was changed that much. We checked the EOM with impedance kit. It has 14.75 MHz peak and the crystal looks nice, so we use it anyway.

2012_01_19.jpg

 The error signal looks nice after cable length adjustment.

I locked the cavity with fast feedback only to measure the transmitted power through the cavity. P_side band is 0.16mW, P_carrier is 0.57mW. So Psideband/Pcarrier ~0.3, this corresponds to modulation depth ~ 0.95. This is close to the calculation Frank did.

I have not tried to measure noise at the error point yet, since I have to flip the phase by 180 for feedback to EOM (TTFSS has a phase flip switch for FAST feedback only). I used a long BNC cable to change the phase by 180 degree, I think making an adaptor for EOM connector to flip the signal might be a better idea to try.

2012_01_19b.jpg

Next step:

  • We will use a Marconi with some amplifiers for LO to drive the EOM and demodulate the signal later. The current function generator is noisy, but it's a good start to see how much better we can gain from resonant EOM.
  • Use Jenne laser to characterize the 14.75 MHz RFPD. Then we can calculate how much gain we get, and how much improvement we need.

 

  788   Thu Jan 19 17:27:35 2012 FrankSummarySafetyStolen Items

More and more stuff got taken from the lab without telling us and now REALLY delay things:

  1. entire Newport lens kit (in wooden box), AR-coating for 1064nm
  2. entire spool of RG405
  3. resonant New Focus EOM, 24.5MHz (resonant PDs and Wenzel OCX are still here)

WHOEVER TOOK IT: BRING IT BACK!!

  787   Thu Jan 19 01:19:31 2012 taraDailyProgressoptic14.75MHz EOM

The 14.75 MHz EOM we have is for visible light, so we went to TNI and borrowed a 14.75 MHz EOM for IR and an 14.75 Mhz resonant RFPD.  I will re-aligned the beam and measure the error signal tomorrow.

 The current function generator can provide power up to 23dBm. So the EOM can be driven around ~ 19 dBm(~2V@ 50Ohm) (-3 dBm for a splitter, -1 dBm for loss in the cable). So we can expect the modulation index to be 0.2*2 = 0.4 rad.

  786   Wed Jan 18 01:23:36 2012 FrankDailyProgressRefCavin-vacuum temp sensors

Finished first pair of temp sensors for cavity radiative heaters. Made some small pieces of copper and soldered the AD590s to it using Indium. Wire is twisted kapton insulated copper wire AWG36. Sensors will be clamped using a single screw size #2 with indium between the sensor plate and surface of the copper tube. Sensors will not be used for temp control, only as monitoring devices. Plan is to sense beat frequency and keep that constant by feeding back to the radiative heater, so no temp sensor is required. Sensors are nice for power estimation and frequency tuning coeff. measurements and heater power limitation if required (which i don't think as the heater resistance is too high to cause any problems even with railing power)

430.jpg432.jpg449.jpg

459.jpg467.jpg

Attachment 4: 458.jpg
458.jpg
  785   Tue Jan 17 21:59:47 2012 taraDailyProgressoptic14.75MHz EOM

As we decided to use lower sideband frequency (14.75MHz, instead of 35.5MHz), I replaced the Broadband EOM with 14.75MHz EOM.

==Motivation==

   The current broadband EOM give only small modulation depth (~0.06 rad with maximum power from the LO, seepsl:745) With a resonant EOM, we can get higher modulation depth with the same amount of power.

   Plus, in general, the RFPD's Q will be also higher at lower frequency, so we should get higher gain to suppress more frequency noise (the exact number of Q has not been measured yet).

==To Do/ Problems==

  We no longer use LIGO's old LO cards. All of the spares in the lab are also broken. We will use a function generator and adjust the cable length to change phase between LO and PD. 

 After I added the resonant EOM to the setup the beam path changed quite a lot, I need to re-aligned the beam before I can see the error signal and lock the cavity.

  784   Fri Jan 13 16:51:29 2012 TaraNotesopticIsolator base, EOM base's height

 Note: new EOM base

This one will have total height of 1.31" . THe height of the EOM (base to aperture) is 0.56". The height of the 4-axis stage (new focus 9071) is 1.06 - 1.18 " (min to max), I use 1.13" as operating height. So the total height is 1.31 + 0.56+ 1.13 = 3" .

 

The drawings of EOM and 4-axis stage can be found here:

EOM

4-axis stage

 

 

Mon Jan 16 17:10:25 2012

Base for Dual periscope is added. This will allow us to mount the plate to the table with screws. Clamps can be used to provide additional support as well.

dual_periscope_base.PDF

Attachment 1: eom_base_2011_01_13.PDF
eom_base_2011_01_13.PDF
  783   Thu Jan 12 16:04:05 2012 TaraSummaryPMCIsolator base, EOM base's height

As mention in the previous corresponding entry, the height of the base for the faraday isolator is not correct. I removed the thing from the table and measure its height again in order to have it fixed. It has to be cut by 3/128 inch from the base (see figure below). The groove for the isolator is well leveled. I got the same height for both ends. I'll bring it back to machine shop tomorrow.

faraday_base.jpg

There are other mechanical parts I need to fix:

  • Broadband EOM's base: Frank noticed that it is a bit too high, and the 4- axis stage height has to be adjusted close to minimum which is not good for mechanical stability.
  • New design for EOM base: If we decide to go for two EOMs for 1)adding side band, and 2) feedback, we need some space for installing the second extra EOM. With the current EOM base design, it takes up to much space along the beam path, so I might need to make a new EOM base.
  • RFPDs' base: Now they are only a plastic post. We will need a sturdier design.
  • Base for dual periscope

I'll try to have these parts before we open the chamber to change the seismic stage.

  782   Wed Jan 11 20:50:08 2012 taraDailyProgressNoiseBudgetcurrent TTFSS noise budget

 As Frank proposed to use side band frequency at 21.5 or 14.75 MHz, I checked the HOM of both frequencies.

See more details about HOM from psl:106

21_5MHz.png

Fig 1: HOM from 21.5 MHz sideband.  At 17th HOM the frequency difference between the sideband and the HOM is ~2MHz, which is much bigger than the cavity's linewidth.

 

14_75MHz.png

Fig2: HOM from 14.75 MHz sideband. The difference between 10th HOM and the resonant frequency is 2.2 MHz, which is much bigger than the cavity linewidth (~75 kHz).

 

It is unlikely that one of the HOMs will resonate inside the cavity if the frequency difference is much larger than the cavity linewidth. So, there is no apparent overlap  we have to worry about for both frequencies.

Thu Jan 12 10:53:29 2012: The code for HOM is on svn.

 

  781   Sat Jan 7 00:45:47 2012 FrankDailyProgressNoiseBudgetcurrent TTFSS noise budget

analyzed the data taken yesterday with the old TTFSS setup. Measured the EP noise (out1, common path, right after the mixer) and converted into frequency noise using an average number of the error signal slope of ~1MHz/V (which by the way is pretty bad).  Plot of electronic noise (demodulated PD signal) shows a noise level of ~20nV/rtHz, not super good but possible. With terminated input (~7nV/rtHz) it is slightly above the noise floor of the analyzer (6nV/rtHz)  which means it's about 4nV/rtHz which is OK.

TTFSS_enoise.png

If we compare the projected noise with the coating thermal noise we see that we are not able to measure anything useful with the current scheme. We need to gain about a factor of 10 to 20 more sensitivity. I suggest we switch to a resonant EOM to increase the modulation index and move to a lower frequency at the same time to increase the sensitivity of our RF PDs (old or new). This might also help us later when we need to stabilize the EOM temperature to minimize the RF-AM. HV-modulation might actually cause RF-AM which we are not able to cancel with our feedback from RF-AM sensing to EOM temp or DC actuator. We should check if we can use existing equipment like a 14.75MHz EOM we have or we can switch to the 21.5MHz which we used before. Switching can be done within one afternoon, two at most.

TTFSS_fnoise.png

 

We've moved on today towards a real TTFSS. Tara made the power cables and we quickly measured the noise of the EP of the new setup. Almost all harmonics are gone. Only tiny pickup of line harmonics is visible. Will make plots later.

  780   Fri Jan 6 12:18:08 2012 taraDailyProgressNoiseBudgetThermoelastic noise from substrate

The correction for thermoelastic noise (by Cerdonio) has been added in the noise budget code. The substrate TE noise estimation from previous code uses Braginsky's result which does not take the spot size of the beam into account.  The code is now on svn.

The TE noise correction given by Cerdonio has to be integrated twice at each frequency point.

cerdonio.png.

The integration limits are 0 to infinity. I use Riemann sum to do the integration. Luckily, the integral converges quickly, so with integration limit from 0 to 10, stepsize = 0.1, the error between Cerdonio's and Braginsky's result is only 1.5% (at high frequency) and the calculation is not exhaustive. This can be changed later if we need better accuracy.

ctn_nb.png

For our experiment the thermoelastic noise is below coating Brownian noise at all frequency.

Attachment 1: TE_noise_compare.png
TE_noise_compare.png
  779   Fri Jan 6 00:53:03 2012 FrankDailyProgressSeismicconical springs

cleaned the conical springs Rana bought. Will characterize and try them tomorrow

Line   Description Ordered Ships Unit Price Total
1 1692K17 Type 302 Stainless Steel Conical Compression Spring, .375" L, .48" L OD, .187" Small OD, .035" Wire Diameter, packs of 3 2
packs
in the morning 14.88
per pack
29.76
2 1692K11 Type 302 Stainless Steel Conical Compression Spring, .250" L, .36" L OD, .125" Small OD, .029" Wire Diameter, packs of 3 2
packs
in the morning 13.52
per pack
27.04
3 1692K21 Type 302 Stainless Steel Conical Compression Spring, .500" L, .72" L OD, .281" Small OD, .055" Wire Diameter, packs of 1 8
packs
in the morning 7.41
per pack
59.28

 

P1820414.JPG

  778   Fri Jan 6 00:26:16 2012 FrankDailyProgressFSSTTFSS equipment moved

FSS is finally table-top equipment, so really a TTFSS . Couldn't find a nice place where the cables don't reach into a beam path and having access to the front panel at the same time. So i added a small breadboard as a mini shelf. Will replace that by a simple plate later. Paths to EOM and RF-PD are now very short. 1U control box will have it's temporary place below the table for now. Same for the power supplies. Will be moved to the rack later when everything is working. Why the hell does this box needs 6 voltages by the way? That's crazy

Below some pictures:

P1820415.JPGP1820425.JPG

 

  777   Wed Jan 4 22:50:59 2012 FrankNotesFSSTTFSS noise hunting started

started noise hunting within the RCAV loop (using the TTFSS electronics) to figure out why the high frequency noise level is higher as we expect.

Started checking the RF-PD. Noise level is still OK, about the same as measured by Raphael during the summer (#661). So the PD is OK for now, not very good but not broken.

Next i checked the noise of the TTFSS input stage (D0901894-A) (mixer etc., used front panel connectors OUT1 and OUT2 in common path, D040105-C).

RF components to estimate mixer LO power:

  • ADT 1-1 (datasheet): isolation transformer, 1:1, loss ~1dB
  • GVA-81 (datasheet): Amplifier, ~10dB gain
  • ADC 10-4 (datasheet): directional coupler, mainline loss ~0.8dB
  • JMS-1H (datasheet): level 17 mixer

--> using an LO input power of 7.7dBm we have ~16dBm at the mixer.

 

Measured the noise level with PD input terminated and with PD connected (dark, without any light). LO power was 7.7dBm @ 35.5MHz for both. Measurement with terminated input was limited by input noise of the SR785, so we need to re-measure it with a preamp, but this already indicates that everything is OK. Noise level with PD is above the instrument noise.

- add plots -
 

  776   Thu Dec 29 23:57:28 2011 FrankPhotosRefCavradiative heater prototype

564.jpg

569.jpg

572.jpg

574.jpg

  775   Thu Dec 29 17:56:39 2011 FrankDailyProgressRefCavradiative heater

went to Downs to get the right crimp tool for the d-sub connectors. It's a DMC crimper for D-sub contacts for UHV connectors (crimp tool). Turns out that the DMC tool does not properly fit the contacts we have, (female pins).. The contacts we have are NASA approved, space qualified, low outgassing crimp contacts (link) and are the same as delivered with the aLIGO, in-vacuum PEEK D-sub connectors. The crimp location seems to be too high for those contacts, but is still working. The insert is the correct one (description), (pdf).

562.jpg

Interesting is that the pins which come with the UHV connector have little colored markings on them as can be seen on the picture (see above link to female pins). Marking can not be removed with methanol, acetone etc.. My guess is that it's OK and actually might be not simply paint.

Will use d-sub contacts to crimp on the nicrome wire and run kapton insulated copper wire to the main vacuum feedthrough. Having the contacts right next to the copper tube makes it a lot easier to work on the whole cavity assembly.

  • wire can be stripped using an old rotary sanding tool where only the matrix is left (no abrasive stuff left).
  • Crimp tool has to be set to setting 1(!) to crimp correctly, default is 5.
  • clamped the wire using #2-56 screws and washers facing round side down to not have sharp edges which could cut the insulation of the wire.
  • added tiny amounts of the super thin kapton tape to further protect the wire
  • total amount of turns is 21
  • total resistance is 154ohms - calculated value was 147ohms
  • crimped femaile contacts to both ends
  774   Thu Dec 29 15:13:08 2011 FrankNotesVacuumchamber weight

Need a rough number for suspension calculations. Used my scale from home which is probably not very accurate.

chamber, including windows, valve and legs : ~41kg (measured)
ion pump, including magnet and insulating CF mount: ~8kg (measured)
stack, including both cavities, copper shields etc: ~14.5kg (from CAD model)

Total : 63.5kg

  773   Thu Dec 29 01:11:32 2011 FrankDailyProgressSeismicAccelerometer noise floor update

found the old data from the seismometer measurements we did some time ago. Measurements agree more or less. Seismic noise floor measured some time ago was higher above 30Hz, but that was measured at the time where we had the construction going on and not during vacation time.

accelerometer_v2.png

accelerometer_v2.fig

  772   Wed Dec 28 01:36:11 2011 FrankDailyProgressRefCavPTFE parts baked. now baking Nichrome wire

removed the PTFE parts for mounting the copper shields from the baking oven.
Now baking the Nichrome wire wrapped onto one of the aluminum radiation shields from the cryo experiment.

P1810539.JPG

  771   Wed Dec 28 01:30:46 2011 FrankDailyProgressSeismicAccelerometer noise floor and table seismic noise

Measured the noise floor of the accelerometer to compare it with the seismic noise we measure with it (and the TFs i want to measure). Suspended it from the table using a very soft spring to make it quiet enough. Resonance frequency of suspension was ~1Hz (small peak visible in spectrum, damping of the thing took forever).

P1810534.JPG

Seismic noise of the table is hitting the noise floor of the accelerometer at about 10Hz. Couldn't find the raw data from the seismometer measurements. Will update this later.

accelerometer.png

  770   Tue Dec 27 19:49:27 2011 FrankNotesopticLens Kit missing!

The entire lens kit (Newport wooden box, v-coating for 1064nm) is missing . Checked other labs but can't find it.

Can't continue work without it

SO PLZ RETURN IT !

ASAP!

  769   Tue Dec 27 16:11:07 2011 FrankNotesSeismicstack TF updated

updated the stack TF using more accurate numbers for the RTV springs and masses and also updated the proposed TF for the stack using the modified RTV springs (see elog #762) (instead of SS springs). The updated model can be found on the svn.

Drilling the hole in the center of the existing springs reduces k to 19150N/m (RTV615 block, ~35Hz resonance (max value measured ~35Hz, minimum ~33Hz) with 0.396kg mass) from the original k=28900N/m (RTV615 block ~43Hz, resonance with 0.396kg mass). It should also be possible to reduce the number of springs from 8 to 4, at least for the top stack. For the bottom we have to check how much clearance we will have left between the two stack plates if we use softer springs and reduce the amount by a factor of two.

 estimated stack TF using measured RTV spring values

stack_TF.png

Ring-down measurement of modified RTV spring

P1810299.JPG

 

  768   Fri Dec 23 18:39:10 2011 FrankNotesRefCavradiative heater

Situation: The heater wire should not go all the way to the vacuum feedthrough connector. Instead the connection from the heater to the connector should be made with some low conductivity material to only add heat where we want it.

So i tried to solder copper wires to the end of the Nichrome wire and failed. Whatever people suggest on the web does not work properly. i tried 6 different kind of tin with different fluxes, 3 separate fluxes incl rosin and acid flux for brazing steel. Oxidized, scratched, etched the surface - nothing worked. you can get a blob of tin at the end but not a clean solder joint.

will go for crimped connections unless someone has an idea....

  767   Fri Dec 23 12:00:54 2011 FrankNotesComputersslow (laser temp) feedback script fixed.

Software feedback to laser temp didn't work correctly. Fixed it.

Script is located in /usr1/epics/psl/scripts

  766   Fri Dec 23 01:11:11 2011 Koji, FrankHowToSeismicKistler Accelerometer preamp Type 5124A1 - gain settings

we were wondering what the gain in the accelerometer readout box is (Type 5124A1). According to the manual the gain can be changed between 0.1 and 100 (manual) but the device is not labeled. So we opened it to check the inside. The gain can be set using DIP switches. Each switch has 4 individual switches, 2 per amplifier stage. There is no label indicating which setting corresponds to which gain. We tried all four combinations while watching an excitation with fixed level and frequency on the FFT analyzer. Turns out that all channels in our preamp are configured to a gain of 1 at the moment. There is also no improvement in SNR when using a higher gain setting.

So the output using our small accelerometers is

~1V/g.

 

P1810511.JPG

  765   Thu Dec 22 20:28:58 2011 FrankDailyProgressRefCavcopper shields

tapped a few #2-56 holes in the top of the copper tube which will be used as a radiative shield and heater for the cavities. The holes are needed to fix the heating wire (AWG32, NiCr80, 11.3Ohms/ft, w/Heavy ML Enamel,  0.0080" Diameter). Polished and cleaned the tube ready for in-vacuum use (not baked, does not fit in chamber). Teflon pieces and screws for mounting the thing got cleaned and are currently vacuum-baked at 80degC over night.
Here a list of thoughts/facts:

  • Wire resistance is about pi*2.5in/12*11.3 = 7.4Ohms/turn.
  • The total heating power for the chamber to bring it from RT to 35degC is about 7W.
  • I would like to drive the heater directly with an low-noise opamp to not have to build a low-noise, high power driver
  • Using standard opamps the maximum voltage will be somewhere around 12V.
  • Assuming 1W of power for each heater is more than plenty we need to go for 20 turns.
  • 20 turns would get us 148 Ohms, so maximum current @12V would be 80mA
  • LT1028/1128 can only drive 25mA, AD797 can directly drive 50mA
  • 50mA and 20 turns would allow (only) 370mW of heating
  • if we go to 25 turns we will have 185Ohms, giving us 465mW of heating power using 50mA @9.25V
  • The whole shield is 10 inches long (the cavity is 8 inches). 
  • A uniform distribution is different as the shield has two cut-outs for the cavity supports.
  • The frequency difference of both cavities at 35C is ~160MHz
  • tuning factor for the cavity is ~155MHz/K (or 6.4mK/MHz)
  • required temp difference will be 1K if we use both AOMs
  • 1K temp gradient requires 22 mW heating power for polished copper cylinder in polished SS cylinder (no conductivity assumed, is wrong but have to measure contact area between PTFE mounts and top stack plate)
  • conductive PTFE pieces are 0.188" thick, assuming an touching area of 0.5"x0.125" this results in a conductive heat transfer of 2.2mW per mount, so a total of 6.6mW for all three mounts
  • --> total heat transfer assuming 1K temperature difference : ~30 mW

 

picture shows one of the shields with some copper wire for test-wrapping

 P1810530.JPG

  764   Thu Dec 22 16:28:10 2011 FrankDailyProgressSeismicre-measuring seismic coupling to cavities

trying to verify our FE model for both cavities. In order to measure the coupling i tried different versions of attaching a "shaker" or PZT  to the table to excite the table in vertical direction. As i'm worried about attaching something to the floating table as it could easily destroy the attached PZT or shaker i try to excite the non-floating table. The best solution i found is to put a PZT right below the vacuum chamber underneath the optical table. It does not look very sturdy but the performance is very good. Driving the PZT with only a few volts using our HV piezo amplifier (old, modified PMC servo) i can easily get an SNR of 100 above the seismic noise. And it's making a lot of noise if you drive it even harder. Turns out that the optical table is a good loudspeaker. Also driving it a low frequencies one can feel the whole lab vibrating!

P1810521.JPG

Measurements of the coupling are currently performed. Check back later...

  763   Sun Dec 11 17:43:09 2011 ranaNotesComputersWD green drives suck

Quote:

again (for the third time now) one of the Western Digital Caviar Green WD15EARS 1.5TB disks failed. We already had two failing in the ATF, now the one in this computer failed too. (WDC WD15EARS)

PLZ DO NOT USE THOSE DRIVES IN FRAMEBUILDERS IN THE FUTURE. THEY WILL CAUSE ONLY TROUBLE.

 Google's drive failure report: SMART is not so smart.

  762   Sat Dec 10 13:13:07 2011 FrankDailyProgressSeismicRTV spring modification

As we know that the spring constant of the current  RTV springs of our seismic stack is too high i tried to find a way to make it softer. The idea is to cut a hole into the center of the spring to remove let's say 50% of the material but keep the existing height and diameter/shape.

The material is very brittle so cutting  them is not easy. I tried one of the cork borers from the 40m. As most of them are f**** i gave the one a wanted to use to the machine shop to fix it. So i have a nice sharp tool for cutting them. In order to center the tool on the spring i've assembled a cutting frame using standard optomechanics (see pictures) and modified the drill diameters of those parts to make everything fit.

The result looks very promising. The first half of the spring is cut very nicely. There i used just a lot of force to cut through the material (straight, no rotation of the tool). On the bottom half i tried to rotate the tool to make the cutting much easier but what happens is that the material breaks into a lot of tiny little pieces which also happens partly to the spring we want as well (see center/bottom of spring). Nevertheless i have a nice tire-shaped RTV spring which i can now measure the spring constant and compare it's performance with the real stainless springs which i have as an alternative. It would be nice to use the RTV as it includes damping in a single device. I'm worried about building a high-Q oscillator using the steel springs without adding additional damping which we currently don't know how to do.

 P1810236.JPGP1810240.JPG

  761   Fri Dec 9 11:55:21 2011 FrankNotesDAQnew channels from second VME crate

Added the following channels. Data is valid from now on.

# RF-AM monitor channels (not used at the moment, no calibration)
[C3:PSL-FSS_RFAM_RCAV]   # monitor for RCAV
[C3:PSL-FSS_RFAM_ACAV]   # monitor for ACAV

# EOM temp stabilization channels
[C3:PSL-FSS_EOM_TSET]  # EOM temp setpoint value (no calibration,units=Volt)
[C3:PSL-FSS_EOM_TACT]  # actual (in-loop) temperature (no calibration, units=Volt)
[C3:PSL-FSS_EOM_IMON]  # EOM heater current monitor - does not exist yet
[C3:PSL-FSS_EOM_SETTEMP]  # Analog output to set EOM temperature - not used at the moment

# aux channels (no calibration, input range -10V to 10V
[C3:PSL-GEN_DAQ1]
[C3:PSL-GEN_DAQ2]
[C3:PSL-GEN_DAQ3]
[C3:PSL-GEN_DAQ4]
[C3:PSL-GEN_DAQ5]
[C3:PSL-GEN_DAQ6]
[C3:PSL-GEN_DAQ7]
[C3:PSL-GEN_DAQ8]

  760   Fri Dec 9 11:49:17 2011 FrankSummaryDAQdaqd problems (which turn out to be not critical)

Talked to Alex about the daqd and it's strange behavior on our machine.

  1. The first thing is that the daqd did not show up as a process even it was running. According to Alex this happens because we only write about 100 EPICS channels so they cpu load is so small that it does not show up all the time so you can easily miss it. The safest way to find out if it's running is to try to telnet into it on port 8087 and check the status.
  2. We have the following error message in our log file (since a long time ago):
    A call to "assert (pE == &entry)" failed in ../../cds/project/daq/fb2/exServer.h line 346
    EPICS Release EPICS R3.14.10- $R3-14-10-RC2$ $2008/10/10 15:01:51$.
    E-mail this message to the author or to tech-talk@aps.anl.gov
    Calling epicsThreadSuspendSelf()

    This happens because it's trying to start an epics ioc using a wrong version of the epics library. We tried different version but couldn't fix it. Looks like the easiest solution would be to re-compile the daqd.
  3. Error message: failed to rename file; errno 2
    This means it can't rename the current frame file it is writing to which can have several reasons. Most often it does not have the rights to do so and a second option is that if the daqd is already running but invisible and you try to start (another) one there is no exception catching for that case and those two processes create a conflict because they try to create/write to the same file.
  4. If you want to grab data using the dataviewer you get the following message even if data exists
    read(); errno=9
    read(); errno=9
    T0= some time and length of data
    No data output

    Unknown problem. Even the data exists and can be grabbed using different command line tools to directly access the frames the dataviewer does not get data. A second weird behavior is that if you request let's say 60 seconds it request 120 seconds. Should not be like that but that's a dataviewer problem.
  759   Fri Dec 9 06:32:24 2011 TaraSummaryPMCrealignment

The V-block's height is a bit too high. The beam height is very close to 3".

Quote:

thanks for the info.

I only did the PMC so far as i had to fix the daq and add the new channels. What's the problem with beam height of the isolator? Is the beam too low or the mount too high? Do you know?

 

 

 

  758   Thu Dec 8 23:56:10 2011 FrankNotesComputersErrors: Linux System Errors
List of standard Linux System Errors
#define EPERM            1      /* Operation not permitted */
#define ENOENT           2      /* No such file or directory */
#define ESRCH            3      /* No such process */
#define EINTR            4      /* Interrupted system call */
#define EIO              5      /* I/O error */
#define ENXIO            6      /* No such device or address */
#define E2BIG            7      /* Arg list too long */
#define ENOEXEC          8      /* Exec format error */
#define EBADF            9      /* Bad file number */
#define ECHILD          10      /* No child processes */
#define EAGAIN          11      /* Try again */
#define ENOMEM          12      /* Out of memory */
#define EACCES          13      /* Permission denied */
#define EFAULT          14      /* Bad address */
#define ENOTBLK         15      /* Block device required */
#define EBUSY           16      /* Device or resource busy */
#define EEXIST          17      /* File exists */
#define EXDEV           18      /* Cross-device link */
#define ENODEV          19      /* No such device */
#define ENOTDIR         20      /* Not a directory */
#define EISDIR          21      /* Is a directory */
#define EINVAL          22      /* Invalid argument */
#define ENFILE          23      /* File table overflow */
#define EMFILE          24      /* Too many open files */
#define ENOTTY          25      /* Not a typewriter */
#define ETXTBSY         26      /* Text file busy */
#define EFBIG           27      /* File too large */
#define ENOSPC          28      /* No space left on device */
#define ESPIPE          29      /* Illegal seek */
#define EROFS           30      /* Read-only file system */
#define EMLINK          31      /* Too many links */
#define EPIPE           32      /* Broken pipe */
#define EDOM            33      /* Math argument out of domain of func */
#define ERANGE          34      /* Math result not representable */
#define EDEADLK         35      /* Resource deadlock would occur */
#define ENAMETOOLONG    36      /* File name too long */
#define ENOLCK          37      /* No record locks available */
#define ENOSYS          38      /* Function not implemented */
#define ENOTEMPTY       39      /* Directory not empty */
#define ELOOP           40      /* Too many symbolic links encountered */
#define EWOULDBLOCK     EAGAIN  /* Operation would block */
#define ENOMSG          42      /* No message of desired type */
#define EIDRM           43      /* Identifier removed */
#define ECHRNG          44      /* Channel number out of range */
#define EL2NSYNC        45      /* Level 2 not synchronized */
#define EL3HLT          46      /* Level 3 halted */
#define EL3RST          47      /* Level 3 reset */
#define ELNRNG          48      /* Link number out of range */
#define EUNATCH         49      /* Protocol driver not attached */
#define ENOCSI          50      /* No CSI structure available */
#define EL2HLT          51      /* Level 2 halted */
#define EBADE           52      /* Invalid exchange */
#define EBADR           53      /* Invalid request descriptor */
#define EXFULL          54      /* Exchange full */
#define ENOANO          55      /* No anode */
#define EBADRQC         56      /* Invalid request code */
#define EBADSLT         57      /* Invalid slot */

#define EDEADLOCK       EDEADLK

#define EBFONT          59      /* Bad font file format */
#define ENOSTR          60      /* Device not a stream */
#define ENODATA         61      /* No data available */
#define ETIME           62      /* Timer expired */
#define ENOSR           63      /* Out of streams resources */
#define ENONET          64      /* Machine is not on the network */
#define ENOPKG          65      /* Package not installed */
#define EREMOTE         66      /* Object is remote */
#define ENOLINK         67      /* Link has been severed */
#define EADV            68      /* Advertise error */
#define ESRMNT          69      /* Srmount error */
#define ECOMM           70      /* Communication error on send */
#define EPROTO          71      /* Protocol error */
#define EMULTIHOP       72      /* Multihop attempted */
#define EDOTDOT         73      /* RFS specific error */
#define EBADMSG         74      /* Not a data message */
#define EOVERFLOW       75      /* Value too large for defined data type */
#define ENOTUNIQ        76      /* Name not unique on network */
#define EBADFD          77      /* File descriptor in bad state */
#define EREMCHG         78      /* Remote address changed */
#define ELIBACC         79      /* Can not access a needed shared library */
#define ELIBBAD         80      /* Accessing a corrupted shared library */
#define ELIBSCN         81      /* .lib section in a.out corrupted */
#define ELIBMAX         82      /* Attempting to link in too many shared libraries */
#define ELIBEXEC        83      /* Cannot exec a shared library directly */
#define EILSEQ          84      /* Illegal byte sequence */
#define ERESTART        85      /* Interrupted system call should be restarted */
#define ESTRPIPE        86      /* Streams pipe error */
#define EUSERS          87      /* Too many users */
#define ENOTSOCK        88      /* Socket operation on non-socket */
#define EDESTADDRREQ    89      /* Destination address required */
#define EMSGSIZE        90      /* Message too long */
#define EPROTOTYPE      91      /* Protocol wrong type for socket */
#define ENOPROTOOPT     92      /* Protocol not available */
#define EPROTONOSUPPORT 93      /* Protocol not supported */
#define ESOCKTNOSUPPORT 94      /* Socket type not supported */
#define EOPNOTSUPP      95      /* Operation not supported on transport endpoint */
#define EPFNOSUPPORT    96      /* Protocol family not supported */
#define EAFNOSUPPORT    97      /* Address family not supported by protocol */
#define EADDRINUSE      98      /* Address already in use */
#define EADDRNOTAVAIL   99      /* Cannot assign requested address */
#define ENETDOWN        100     /* Network is down */
#define ENETUNREACH     101     /* Network is unreachable */
#define ENETRESET       102     /* Network dropped connection because of reset */
#define ECONNABORTED    103     /* Software caused connection abort */
#define ECONNRESET      104     /* Connection reset by peer */
#define ENOBUFS         105     /* No buffer space available */
#define EISCONN         106     /* Transport endpoint is already connected */
#define ENOTCONN        107     /* Transport endpoint is not connected */
#define ESHUTDOWN       108     /* Cannot send after transport endpoint shutdown */
#define ETOOMANYREFS    109     /* Too many references: cannot splice */
#define ETIMEDOUT       110     /* Connection timed out */
#define ECONNREFUSED    111     /* Connection refused */
#define EHOSTDOWN       112     /* Host is down */
#define EHOSTUNREACH    113     /* No route to host */
#define EALREADY        114     /* Operation already in progress */
#define EINPROGRESS     115     /* Operation now in progress */
#define ESTALE          116     /* Stale NFS file handle */
#define EUCLEAN         117     /* Structure needs cleaning */
#define ENOTNAM         118     /* Not a XENIX named type file */
#define ENAVAIL         119     /* No XENIX semaphores available */
#define EISNAM          120     /* Is a named type file */
#define EREMOTEIO       121     /* Remote I/O error */
#define EDQUOT          122     /* Quota exceeded */

#define ENOMEDIUM       123     /* No medium found */
#define EMEDIUMTYPE     124     /* Wrong medium type */
  757   Thu Dec 8 23:54:05 2011 FrankNotesComputersdaq still broken, but different

checked the source code for the daqd. There is only one line where one of the messages occurs. Couldn't find the part of the source code for the second error message

/* Write out a frame */
        int nwritten = write (fd, ost -> str (), image_size);
        if (nwritten == image_size) {
          if (rename(_tmpf, tmpf)) {
            system_log(1, "failed to rename file; errno %d", errno);
            fsd.report_lost_frame ();
            set_fault ();
          } else {
            DEBUG(3, cerr << "frame " << frame_cntr << "(" << frame_number << ") is written out" << endl);
            // Successful frame write
            fsd.update_dir (gps, gps_n, frame_file_length_seconds, dir_num);
          }
        } else {
          system_log(1, "failed to write full frame out; errno %d", errno);
          fsd.report_lost_frame ();
          set_fault ();
        }
        close (fd);
        TNF_PROBE_0(daqc_c_framer_frame_write_end, "daqd_c::framer", "frame write");

looks like the daqd has problems writing the frames to disk. Standard Error #  2: No such file or directory

Quote:

replaced the hard drive, created an new ext3 file system and mounted it as /frames/full. After forced fsck for the other drives computer is back online. Daqd is working in principle, but  it crashes after a minute or so.

Most often error message:   [Thu Dec  8 20:30:41 2011] failed to rename file; errno 2

Error message changes from time to time: [Thu Dec  8 20:09:45 2011] framer(): failed to rename file; errno 2

I don't know which file and why does the error message change?. Have to contact Alex...

 

  756   Thu Dec 8 21:31:26 2011 FrankNotesComputersdaq still broken, but different

replaced the hard drive, created an new ext3 file system and mounted it as /frames/full. After forced fsck for the other drives computer is back online. Daqd is working in principle, but  it crashes after a minute or so.

Most often error message:   [Thu Dec  8 20:30:41 2011] failed to rename file; errno 2

Error message changes from time to time: [Thu Dec  8 20:09:45 2011] framer(): failed to rename file; errno 2

I don't know which file and why does the error message change?. Have to contact Alex...

  755   Thu Dec 8 21:06:10 2011 FrankSummaryPMCrealignment

thanks, for the info.

I only did the PMC so far as i has to fix the daq and add the new channels. What's the problem with beam height of the isolator? Is the beam too low or the mount too high? Do you know?

Quote:

A reminder for Frank about the setup, 

  • The AOM's case on RCAV beam path is not screwed down to the body, so it will be blocking the beam.
  • The height for Faraday isolator behind the PMC is not correct, i think only ~80% of the power is transmitted. I haven't had it fixed yet
  • When I measured the visibility of RCAV, I scanned the cavity and minimize the dip as seen on the REFL RFPD. I got sth ~90% transmission, but when I measured the actual tranmitted power, the visibility is lower than that, maybe only~ 75%.
  • There is some weird reflection on ACAV RFPD, I'm not sure if it comes from the window or not.

 

  754   Thu Dec 8 20:28:06 2011 taraSummaryPMCrealignment

A reminder for Frank about the setup, 

  • The AOM's case on RCAV beam path is not screwed down to the body, so it will be blocking the beam.
  • The height for Faraday isolator behind the PMC is not correct, i think only ~80% of the power is transmitted. I haven't had it fixed yet
  • When I measured the visibility of RCAV, I scanned the cavity and minimize the dip as seen on the REFL RFPD. I got sth ~90% transmission, but when I measured the actual tranmitted power, the visibility is lower than that, maybe only~ 75%.
  • There is some weird reflection on ACAV RFPD, I'm not sure if it comes from the window or not.
  753   Thu Dec 8 15:59:21 2011 FrankSummaryPMCrealignment

started realigning everything from scratch and calibrate all channels right. PMC optical power channels need re-calibration.

input beam to PMC: 28.9mW
PMC transmitted beam: 23.4mw
transmission trough curved mirror: 1.63mW

--> only 81% visibility (87% including back mirror leakage), but should be enough for what we want to do. Don't know the loss mechanism (didn't investigate)

  752   Thu Dec 8 15:03:10 2011 FrankNotesComputersWD green drives suck

again (for the third time now) one of the Western Digital Caviar Green WD15EARS 1.5TB disks failed. We already had two failing in the ATF, now the one in this computer failed too. (WDC WD15EARS)

PLZ DO NOT USE THOSE DRIVES IN FRAMEBUILDERS IN THE FUTURE. THEY WILL CAUSE ONLY TROUBLE.

ELOG V3.1.3-